OpenStudy (anonymous):

Mike went to a store and spent half his money. Now he has as many pennies as he had dollars before, but only half as many dollars as he had pennies before. How much did he spent?

OpenStudy (anonymous):

*spend

OpenStudy (anonymous):

(100)d + p = m and (1/2)(100)p + d = (1/2)m multiply the second equation by 2 and set the left sides of both equations equal to each other. Establish a ratio of d/p and you will get the original amount of dollars and pennies and from that original amount of money you can divide by 2.

OpenStudy (anonymous):

Am I supposed to solve the frat equation for p and plug into the second equation?

OpenStudy (anonymous):

First* not frat

OpenStudy (anonymous):

This type of problem is a bit different from the usual "2 equations in 2 variables" because you have a third variable "m" for money. So, this type of problem has to use proportions. It won't be that hard. Once you follow the directions in the first post exactly, you will get one equation in 2 variables (d and p). That's why you have to use a proportion. Work half of the first post out, and if you get that far and are still stuck, I'll help you finish it.

OpenStudy (anonymous):

Start by multiplying the second equation (both sides) by 2.

OpenStudy (anonymous):

If you want help on this problem, you'll have to come back because I'm going to have to leave soon, @Skip2mylou426

OpenStudy (anonymous):

So if I multiply the second equation by 2, I get 100p+2d=m

OpenStudy (anonymous):

That's good. Now set the left side of that equal to the left side of equation #1. That is permissible because both equal "m".

OpenStudy (anonymous):

Ok, so 99p=98d

OpenStudy (anonymous):

Very good. Establish a ratio of d to p. I'd use the fraction d/p equal to something, from what you just typed.

OpenStudy (anonymous):

d/p=99/98

OpenStudy (anonymous):

Very good. What we have is of course a ratio. But we can use that ratio as actual values of d and p. In other words, we can take as our original # of dollars to be 99 and our original # of pennies to be 98. That gives use $99.98 to start with. Notice what happens when you divide that by 2.

OpenStudy (anonymous):

So, $99.98 divided by 2 is $49.99 Notice the switch of 99.

OpenStudy (anonymous):

And 49 is half of 98. And that fulfills the requirements of the problem.

OpenStudy (anonymous):

Gotcha! Thank you for your help!

OpenStudy (anonymous):

This was actually a very fun problem. That's if you like this type of fun! uw!